OpenStudy (anonymous):

find the limit of the trigonometric function. lim as x approaches 0 sec2x

OpenStudy (anonymous):

just plug in 0 for x, you get sec(0) = 1

OpenStudy (anonymous):

I bet its really easy but I don't see why its one

OpenStudy (anonymous):

The value of sec (0) is 1. Why? sec (0) = 1/cos(0) = 1/1 = 1 secant function is the reciprocal of the cosine function.

OpenStudy (anonymous):

Makes sense?

OpenStudy (anonymous):

can that be shown on the unit circle

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you show me

OpenStudy (anonymous):

On the unit circle, a circle whose radius is 1 and center is at (0,0). Look at the point (0,1). The x-coordinate is cos x. So cos 0 = 1. Since cos 0 = 1, then sec 0 = 1 becuase secant and cosecant are reciprocal functions.

OpenStudy (anonymous):

I think I kind of see it...so lets say the radius for the unit circle was two instead of 1. the sec of zero would be 2

OpenStudy (anonymous):

here u go

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcome.